# NVS-Solver: Video Diffusion Model as Zero-Shot Novel View Synthesizer

### v1

@@ Abstract
By harnessing the potent generative capabilities of pre-trained large video diffusion models, we propose NVS-Solver, a new novel view synthesis (NVS) paradigm that operates without the need for training. NVS-Solver adaptively modulates the diffusion sampling process with the given views to enable the creation of remarkable visual experiences from single or multiple views of static scenes or monocular videos of dynamic scenes. Specifically, built upon our theoretical modeling, we iteratively modulate the score function with the given scene priors represented with warped input views to control the video diffusion process. Moreover, by theoretically exploring the boundary of the estimation error, we achieve the modulation in an adaptive fashion according to the view pose and the number of diffusion steps. Extensive evaluations on both static and dynamic scenes substantiate the significant superiority of our NVS-Solver over state-of-the-art methods both quantitatively and qualitatively. Source code in https://github.com/ZHU-Zhiyu/NVS__\__Solver.
@@ 1 Introduction
In the realm of computer vision and graphics, novel view synthesis (NVS) from limited visual data remains a formidable challenge with profound implications across various domains, from entertainment [46, 16] to autonomous navigation [1, 24] and beyond [3, 62, 35, 30, 20]. However, addressing this challenge demands a sophisticated method capable of extracting meaningful information from sparse visual inputs and synthesizing coherent representations of unseen viewpoints [63, 58]. In this context, the emerging field of deep learning has witnessed remarkable strides, particularly with the advent of advanced generative models [48, 7, 11].
Recently, diffusion models [13, 41, 45] have garnered significant attention due to their exceptional ability to synthesize visual data. A prominent area of focus within this domain is video diffusion models [5, 21, 14, 31, 18, 21, 53], which have gained popularity for their remarkable video generation capabilities.
In this paper, we address the problem of novel view synthesis from single or multiple views of static scenes or monocular videos of dynamic scenes, leveraging the pre-trained large video diffusion model without additional training. Specifically, we theoretically formulate the NVS-oriented diffusion process as guided sampling, in which the intermediate diffusion results are modulated with the scene information from the given views. Moreover, we empirically and theoretically investigate the potential distribution of the error map to achieve adaptive modulation in the reverse diffusion process, with a reduced estimation error boundary.
@@ 2 Related Work
Diffusion Model indicates a kind of deep generative model [39], which is inspired by non-equilibrium statistical physics by iteratively appending noise into image data and then reversely removing noise and transferring to noise-free data distribution. Ho et al. [13] proposed a variance preserving (VP) diffusion denoising probabilistic model via progressively removing noise. Song et al. [45, 43, 44, 42] proposed a score-based image generation model by iteratively calculating the derivative of data distribution and utilizing the stochastic differential equation (SDE) [2] or ordinary differential equation (ODE)-based solvers [27, 45] to reverse the noise-adding process and derive the clean image distribution. Inspired by the success of image diffusion models, many works attempted to build video diffusion model to directly achieve video generation prompted by text [21, 53] or a single image [5, 18].
@@
Diffusion sampling algorithm aims to speed up or control the diffusion process via regularizing or re-directing the reverse trajectory of diffusion models [25, 26, 61, 60, 50, 6, 9]. Song et al. [41] proposed denoising diffusion implicit models (DDIM) to accelerate the diffusion sampling by jumping to the clean image-space at each step. Lu et al. [25, 26, 61] proposed a series of fast ODE diffusion solvers given an analytic solution of ODE by its semi-linear nature. Moreover, the integration of non-linear network-related parts was further approximated by its Taylor series. Zhang et al. [60] explored the huge variance of distribution shift and then decoupled an exponential variance component from the score estimation model, thus reducing the discretization error. Chung et al. [9] proposed to regularize the intermediate derivative from the reconstruction process to achieve image restoration. Wang et al. [50] decoupled the image restoration into range-null spaces and focus on the reconstruction of null space, which contains the degraded information.
@@
Novel view synthesis (NVS) targets at generating images of a scene from viewpoints not presented in the original data and has been a subject of extensive research in computer vision and graphics [33, 8, 47, 3, 36]. Early methods in this domain often relied on geometric approaches, such as multi-view stereo reconstruction [38, 17] and structure-from-motion [37, 32, 32] techniques, to synthesize novel viewpoints from multiple images captured from different angles. The advent of deep learning has revolutionized the field of NVS, enabling the synthesis of realistic images from pre-trained feature volumes. Neural Radiance Fields (NeRF) [30, 34, 4, 28, 23] enabled stunningly detailed renders from 2D images through volumetric rendering. Additionally, differentiable rendering has allowed gradients of rendering outputs with respect to scene parameters, facilitating direct optimization of scene geometries, lighting, and materials. Recently, 3D Gaussian Splatting [20] presented an explicit representation of the scene.
@@
While these methods have shown promising results, they often suffer from limitations such as dependence on dense scene geometry, challenges in handling complex scene dynamics, and being hard to generalize [20]. Moreover, they require significant computational resources and suffer from artifacts such as disocclusions and view-dependent effects.
@@
3 Preliminary
We briefly introduce some preliminary knowledge of diffusion models [39], which facilitates our subsequent analyses. We also refer readers to [43, 45] for more details. Generally, the forward SDE process of the latent diffusion model can be formulated as
@@ 3 Preliminary
where $\bm{\mu}_{t}=\mathcal{X}_{\bm{\theta}}(\bm{x}_{t},t)$ is the estimated clean image from the noised image $\bm{x}_{t}$ at step $t$ by the denosing U-Net $\mathcal{X}_{\bm{\theta}}(\cdot,~{}\cdot)$ [^1] For simplicity, here we combine some parameterizations stemmed from EDM [19] into $\mathcal{X}_{\bm{\theta}}(\cdot)$.. Finally, considering the special case of variance exploding (VE) diffusion process [45] of the stable video diffusion (SVD) [5], Eq. (5) can be further simplified as
@@ 4 Proposed Method
Motivated by the powerful generative capability with realistic and consistent frames by large video diffusion models, we aim to adapt the pre-trained video diffusion model to the task of NVS without any additional training, leading to NVS-Solver. Generally, NVS-Solver focuses on effectively incorporating the prior information of the given scene to guide the reverse sampling of a video diffusion process towards high-quality and desired novel views. Specifically, we first theoretically reformulate NVS-oriented reverse diffusion sampling by modulating the score function with the given views (Sec. 4.1). Based on theoretically exploring the boundary of the diffusion estimation and depth-based warping errors, we propose to adaptively modulate the prior of the given view into diffusion process to reduce the potential boundary of the estimation error (Sec. 4.2).
In the following, we take the single view-based NVS to illustrate our method, as summarized in Algorithm 1, which involves from a given view $\mathbf{X}_{0,\mathbf{p}_{0}}$ with pose $\mathbf{p}_{0}$, reconstructing $N-1$ novel views at target poses $\{\mathbf{p}_{1},\cdots,\mathbf{p}_{i},\cdots,\mathbf{p}_{N-1}\}$, denoted as $\mathbf{X}_{0,\mathbf{p}_{i}}$. Note that our method is suitable for scenarios involving NVS from multiple views, as well as from monocular videos, as explained in Sec. 4.3. Also, we consider the pre-trained image-to-video diffusion model SVD [5].
@@ 4.1 Scene Prior Modulated Reverse Diffusion Sampling
According to the intensity function [29], we can formulate the relationship between $\mathbf{X}_{0,\mathbf{p}_{0}}$ and $\mathbf{X}_{0,\mathbf{p}_{i}}$ through Taylor expansion:
@@
where $\mathcal{I}(\cdot)$ is the intensity function, $\Delta\mathbf{p}:=\mathbf{p}_{i}-\mathbf{p}_{0}$ is the pose variation, and $\mathcal{O}^{2}(\Delta\mathbf{p})$ is the high order Taylor expansions. Based on the depth-driven image warping operation, we further have
Algorithm 1 NVS-Solver for Synthesizing Novel Views from Single Images
@@
where $\Delta\mathbf{D}:=\mathbf{D}-\widehat{\mathbf{D}}$, $\widehat{\mathbf{X}}_{0,\mathbf{p}_{i}}$ refers to the warped view from $\mathbf{p}_{0}$ to pose $\mathbf{p}_{i}$ through $\widehat{\mathbf{D}}$, and $\mathcal{E}_{T}$ is a residual term that contains high-order Taylor expansion series and warping error.
Based on the above analysis that the rendered novel view at pose $\mathbf{p}_{i}$ is highly correlated with $\widehat{\mathbf{X}}_{0,\mathbf{p}_{i}}$ and $\bm{\mu}_{t,\mathbf{p}_{i}}$, an effective score function for NVS should take advantage of them. Then, we formulate the score function of NVS-oriented reverse diffusion sampling, which is modulated with the given scene information as
$\widetilde{\bm{\mu}}_{t,\mathbf{p}_{i}}=\mathop{\mathrm{argmin}}_{\bm{\mu}}\|\bm{\mu}-\bm{\mu}_{t,\mathbf{p}_{i}}\|_{2}^{2}+\lambda(t,\mathbf{p}_{i})\|\bm{\mu}-\widehat{\mathbf{X}}_{0,\mathbf{p}_{i}}\|_{2}^{2},$ (11)
@@
With the optimized clean image expectation term $\widetilde{\bm{\mu}}_{t,\mathbf{p}_{i}}$, we further utilize two ways to guide the reverse sampling of the pre-trained SVD: (1) Directly Guided Sampling, which is fast with limited quality; and (2) Posterior Sampling, which is slow but more effective. Specifically, the former directly replaces the $\bm{\mu}_{t,\mathbf{p}}$ in Eq. (7) with $\widetilde{\bm{\mu}}_{t,\mathbf{p}_{i}}$ as
@@ 4.2 Adaptive Determination of $\lambda(t,\mathbf{p}_{i})$
Although we have reformulated the NVS-oriented diffusion process in the preceding section, the value of $\lambda(t,\mathbf{p}_{i})$ in Eqs. (11) and (12), which is crucial to the quality of synthesized views, has to be appropriately determined. Here, we theoretically explore the formulation of $\lambda(t,\mathbf{p}_{i})$ via analyzing the upper boundary of the estimation error of $\widetilde{\bm{\mu}}_{t,\mathbf{p}_{i}}$.
@@
$\widehat{\lambda}(t,\mathbf{p}_{i})=\mathop{\mathrm{argmin}}_{\lambda(t,\mathbf{p}_{i})}\frac{1}{1+\lambda(t,\mathbf{p}_{i})}\mathbb{E}_{\mathbf{X}_{t}\sim\mathcal{P}(\mathbf{X}_{t})}(\mathcal{E}_{D})+\frac{\lambda(t,\mathbf{p}_{i})}{1+\lambda(t,\mathbf{p}_{i})}\mathbb{E}_{\mathbf{X}\sim\mathcal{P}(\mathbf{X})}(\mathcal{E}_{P})+v_{1}|\log(\lambda(t,\mathbf{p}_{i}))|,$ (16)
where $v_{1}>0$ and the last regularization term prevents the weights from being overfitting on the empirically estimated errors. In the following, we will provide the explicit formulations of the two error terms $\mathcal{E}_{D}=\|\bm{\mu}_{t,\mathbf{p}_{i}}-\widetilde{\bm{\mu}}_{t,\mathbf{p}_{i}}^{*}\|_{2}$ and $\mathcal{E}_{P}=\|\widehat{\mathbf{X}}_{0,\mathbf{p}_{i}}-\widetilde{\bm{\mu}}_{t,\mathbf{p}_{i}}^{*}\|_{2}$  [^2] There is also inherent discretization error as investigated by [25, 26, 60]. However, we could reduce such error terms by enlarging the number of reversing steps., based on theoretical analyses and experimental observations.
Diffusion Estimation Error $\mathcal{E}_{D}$. This error is caused due to the fact that the diffusion model cannot perfectly estimate $\widetilde{\bm{\mu}}_{t,\mathbf{p}_{i}}^{*}$ from the given noised latent $\mathbf{X}_{t}$. Recent works [60, 61] indicate that the derivative $S_{\theta}(\mathbf{X}_{t},t)$ of SDE-based diffusion models varies intensely when $\sigma(t)$ is huge. Accordingly, large values of $\sigma$ would potentially lead to large errors. Moreover, we experimentally investigated the correlation of $\|\mathcal{E}_{D}\|_{2}$ with $\sigma(t)$. As demonstrated in in Fig. 1 (a). $\|\mathcal{E}_{D}\|_{2}$ gradually decreases with the diffusion reverse process (decreasing of $\sigma(t)$). Thus, we empirically formulate $\|\mathcal{E}_{D}\|_{2}=v_{2}\sigma(t)$, where $0<v_{2}$.
@@
$\widehat{\lambda}(t,\mathbf{p}_{i})=\mathop{\mathrm{argmin}}_{\lambda(t,\mathbf{p}_{i})}\frac{v_{2}\sigma(t)}{1+\lambda(t,\mathbf{p}_{i})}+\frac{\lambda(t,\mathbf{p}_{i})v_{3}\|\Delta\mathbf{p}\|_{2}}{1+\lambda(t,\mathbf{p}_{i})}+v_{1}|\log(\lambda(t,\mathbf{p}_{i}))|,$ (17)
whose closed-form solution is

### v2

@@ Abstract
By harnessing the potent generative capabilities of pre-trained large video diffusion models, we propose a new novel view synthesis paradigm that operates without the need for training. The proposed method adaptively modulates the diffusion sampling process with the given views to enable the creation of visually pleasing results from single or multiple views of static scenes or monocular videos of dynamic scenes. Specifically, built upon our theoretical modeling, we iteratively modulate the score function with the given scene priors represented with warped input views to control the video diffusion process. Moreover, by theoretically exploring the boundary of the estimation error, we achieve the modulation in an adaptive fashion according to the view pose and the number of diffusion steps. Extensive evaluations on both static and dynamic scenes substantiate the significant superiority of our method over state-of-the-art methods both quantitatively and qualitatively. The source code can be found on https://github.com/ZHU-Zhiyu/NVS_Solver.
Equal Contribution.*Corresponding Author. This project was supported in part by the NSFC Excellent Young Scientists Fund 62422118, in part by the Hong Kong RGC under Grants 11219422 and 11219324, and in part by Hong Kong UGC under grants UGC/FDS11/E02/22 and UGC/FDS11/E03/24.
[…]
[loop, autoplay, width=]38demo/palace/195
[…]
[loop, autoplay, width=]10demo/kda_dynamic_1/125
[…]
[loop, autoplay, width=]10demo/playground/024
[…]
[width=loop, autoplay]10demo/kda_input_video/125
[…]
[width=loop, autoplay]10demo/camel_input_video/125
[…]
[width=loop, autoplay]2demo/input/playground12
[…]
[width=loop, autoplay]2demo/input/scan3_12
[…]
[loop, autoplay, width=]38demo/camel_ours/195
[…]
[loop, autoplay, width=]10demo/camel_dynamic_2/125
[…]
[loop, autoplay, width=]10demo/scan3/024
Figure 1: Visual demonstrations of the proposed algorithm with input of (a) single-view, (b) monocular video, and (c) multi-view (2 views). The middle row refers to the algorithm’s inputs for each scenario. Please use Adobe Acrobat to display these videos.
@@ 1 Introduction
In the realm of computer vision and graphics, novel view synthesis (NVS) from limited visual data remains a formidable challenge with profound implications across various domains, from entertainment (Tewari et al., 2020; Jiang et al., 2024) to autonomous navigation (Adamkiewicz et al., 2022; Kwon et al., 2023) and beyond (Avidan & Shashua, 1997; Zhou et al., 2016; Riegler & Koltun, 2020; Mildenhall et al., 2021; Kerbl et al., 2023). However, addressing this challenge demands a sophisticated method capable of extracting meaningful information from sparse visual inputs and synthesizing coherent representations of unseen viewpoints (Zou et al., 2024; Zhang et al., 2021). In this context, the emerging field of deep learning has witnessed remarkable strides, particularly with the advent of advanced generative models (Voleti et al., 2024; Chen et al., 2023; Gu et al., 2023).
Recently, diffusion models (Ho et al., 2020; Song et al., 2020; 2021b) have garnered significant attention due to their exceptional ability to synthesize visual data. A prominent area of focus within this domain is video diffusion models (Blattmann et al., 2023; Khachatryan et al., 2023; Ho et al., 2022; Ni et al., 2023; Karras et al., 2023; Khachatryan et al., 2023; Wu et al., 2023), which have gained popularity for their remarkable video generation capabilities. In this paper, we explore the problem of NVS from single or multiple views of static scenes or monocular videos of dynamic scenes, leveraging the pre-trained large video diffusion model without additional training. Specifically, we theoretically formulate the NVS-oriented diffusion process as guided sampling, in which the intermediate diffusion results are modulated with the scene information from the given views. Moreover, we empirically and theoretically investigate the potential distribution of the error map to achieve adaptive modulation in the reverse diffusion process, with a reduced estimation error boundary.
@@ 2 Related Work
Diffusion Model indicates a kind of deep generative model (Sohl-Dickstein et al., 2015), which is inspired by non-equilibrium statistical physics by iteratively appending noise into image data and then reversely removing noise and transferring to noise-free data distribution. (Ho et al., 2020) proposed a variance preserving (VP) diffusion denoising probabilistic model via progressively removing noise. (Song et al., 2021b; Song & Ermon, 2019; 2020; Song et al., 2021a) proposed a score-based image generation model by iteratively calculating the derivative of data distribution and utilizing the stochastic differential equation (SDE) (Anderson, 1982) or ordinary differential equation (ODE)-based solvers (Maoutsa et al., 2020; Song et al., 2021b) to reverse the noise-adding process and derive the clean image distribution. Inspired by the success of image diffusion models, many works attempted to build video diffusion model to directly achieve video generation prompted by text (Khachatryan et al., 2023; Wu et al., 2023) or a single image (Blattmann et al., 2023; Karras et al., 2023).
@@
Diffusion sampling algorithm aims to speed up or control the diffusion process via regularizing or re-directing the reverse trajectory of diffusion models (Lu et al., 2022a; b; Zheng et al., 2024; Zhang & Chen, 2022; Wang et al., 2023b; Cao et al., 2024; Chung et al., 2023; 2022a; 2022b). (Song et al., 2020) proposed denoising diffusion implicit models (DDIM) to accelerate the diffusion sampling by jumping to the clean image-space at each step. (Dhariwal & Nichol, 2021) utilized the classifier to guide the sampling process of diffusion model for controlling the results’ categories. (Lu et al., 2022a; b; Zheng et al., 2024) proposed a series of fast ODE diffusion solvers given an analytic solution of ODE by its semi-linear nature. Moreover, the integration of non-linear network-related parts was further approximated by its Taylor series. (Zhang & Chen, 2022) explored the huge variance of distribution shift and then decoupled an exponential variance component from the score estimation model, thus reducing the discretization error. (Chung et al., 2023) proposed to regularize the intermediate derivative from the reconstruction process to achieve image restoration. (Wang et al., 2023b) decoupled the image restoration into range-null spaces and focused on the reconstruction of null space, which contains the degraded information.
@@
Novel view synthesis (NVS) targets at generating images of a scene from viewpoints not presented in the original data and has been a subject of extensive research in computer vision and graphics (Park et al., 2017; Choi et al., 2019; Tretschk et al., 2021; You et al., 2023; Avidan & Shashua, 1997; Riegler & Koltun, 2021; You & Hou, 2024). Early methods in this domain often relied on geometric approaches, such as multi-view stereo reconstruction (Seitz et al., 2006; Jin et al., 2005) and structure-from-motion (Schonberger & Frahm, 2016; Özyeşil et al., 2017) techniques, to synthesize novel viewpoints from multiple images captured from different angles. The advent of deep learning has revolutionized the field of NVS, enabling the synthesis of realistic images from pre-trained feature volumes. (Rombach et al., 2021; Ren & Wang, 2022) employed autoregressive Transformer to synthesize 3D scene from single image. Neural Radiance Fields (NeRF) (Mildenhall et al., 2021; Pumarola et al., 2021; Barron et al., 2021; Martin-Brualla et al., 2021; Kosiorek et al., 2021) enabled stunningly detailed renders from 2D images through volumetric rendering. Additionally, differentiable rendering has allowed gradients of rendering outputs with respect to scene parameters, facilitating direct optimization of scene geometries, lighting, and materials. Recently, 3D Gaussian Splatting (Kerbl et al., 2023; Liu et al., 2024) presented an explicit representation of the scene.
@@
While these methods have shown promising results, they often suffer from limitations such as dependence on dense scene geometry, challenges in handling complex scene dynamics, and being hard to generalize (Kerbl et al., 2023). Moreover, they require significant computational resources and suffer from artifacts such as disocclusions and view-dependent effects. More recently,  (Charatan et al., 2024; Chen et al., 2025) introduced the use of feed-forward 3D Gaussians for novel view interpolation. (Wu et al., 2024b; Watson et al., 2023; Yu et al., 2023; Cai et al., 2023; Tseng et al., 2023; Chan et al., 2023; Sargent et al., 2024) explored the integration of 3D reconstruction techniques with diffusion models.
@@
3 Preliminary of Diffusion Models
We briefly introduce some preliminary knowledge of diffusion models (Sohl-Dickstein et al., 2015), which facilitates our subsequent analyses. We also refer readers to (Song & Ermon, 2019; Song et al., 2021b) for more details. Generally, the forward SDE process of the latent diffusion model can be formulated as
@@ 3 Preliminary of Diffusion Models
where $\bm{\mu}_{t}=\mathcal{X}_{\bm{\theta}}(\bm{x}_{t},t)$ is the estimated clean image from the noised image $\bm{x}_{t}$ at step $t$ by the denosing U-Net $\mathcal{X}_{\bm{\theta}}(\cdot,~{}\cdot)$ [^1] For simplicity, here we combine some parameterizations stemmed from EDM (Karras et al., 2022) into $\mathcal{X}_{\bm{\theta}}(\cdot)$.. Finally, considering the special case of variance exploding (VE) diffusion process (Song et al., 2021b) of the stable video diffusion (SVD) (Blattmann et al., 2023), Eq. (5) can be further simplified as
@@ 4 Proposed Method
Motivated by the powerful generative capability with realistic and consistent frames by large video diffusion models, we aim to adapt the pre-trained video diffusion model to the task of NVS without any additional training, leading to a score modulation-based approach. Generally, our method warps the input image set to the target view and leverages these warped images as prior information to guide the reverse sampling of a video diffusion process. This not only ensures the accurate generation of the warped regions but also facilitates the meaningful synthesis of unknown or occluded areas, resulting in high-quality and visually consistent novel views with enhanced fidelity and coherence. Specifically, we first theoretically reformulate NVS-oriented reverse diffusion sampling by modulating the score function with the given views (Sec. 4.1). Based on theoretically exploring the boundary of the diffusion estimation and depth-based warping errors, we propose to adaptively modulate the prior of the given view into the diffusion process to reduce the potential boundary of the estimation error (Sec. 4.2).
In the following, we take the single view-based NVS to illustrate our method, as summarized in Algorithm 1, which involves from a given view $\mathbf{X}_{0,\mathbf{p}_{0}}$ with pose $\mathbf{p}_{0}$, reconstructing $N-1$ novel views at target poses $\{\mathbf{p}_{1},\cdots,\mathbf{p}_{i},\cdots,\mathbf{p}_{N-1}\}$, denoted as $\mathbf{X}_{0,\mathbf{p}_{i}}$, . Note that our method is suitable for scenarios involving NVS from multiple views, as well as from monocular videos, as explained in Sec. 4.3. Also, we consider the pre-trained image-to-video diffusion model SVD (Blattmann et al., 2023).
@@ 4.1 Scene Prior Modulated Reverse Diffusion Sampling
According to the intensity function (McMillan & Bishop, 1995), we can formulate the relationship between $\mathbf{X}_{0,\mathbf{p}_{0}}$ and $\mathbf{X}_{0,\mathbf{p}_{i}}$ through Taylor expansion:
@@
where $\mathcal{I}(\cdot)$ is the intensity function, $\Delta\mathbf{p}:=\mathbf{p}_{i}-\mathbf{p}_{0}$ is the pose variation, and $\mathcal{O}^{2}(\Delta\mathbf{p})$ is the high order Taylor expansions. Based on the depth-driven image-warping operation, we further have
Algorithm 1 Zero-shot NVS from Single Images
@@
where $\Delta\mathbf{D}:=\mathbf{D}-\widehat{\mathbf{D}}$, $\widehat{\mathbf{X}}_{0,\mathbf{p}_{i}}$ refers to the warped view from $\mathbf{p}_{0}$ to pose $\mathbf{p}_{i}$ through $\widehat{\mathbf{D}}$, and $\mathcal{E}_{T}$ is a residual term that contains high-order Taylor expansion series and warping error. Based on the above analysis that the rendered novel view at pose $\mathbf{p}_{i}$ is highly correlated with $\widehat{\mathbf{X}}_{0,\mathbf{p}_{i}}$ and $\bm{\mu}_{t,\mathbf{p}_{i}}$, an effective score function for NVS should take advantage of them. Then, we formulate the score function of NVS-oriented reverse diffusion sampling, which is modulated with the given scene information as
$\widetilde{\bm{\mu}}_{t,\mathbf{p}_{i}}=\operatorname*{arg\,min}_{\bm{\mu}}\|\bm{\mu}-\bm{\mu}_{t,\mathbf{p}_{i}}\|_{2}^{2}+\lambda(t,\mathbf{p}_{i})\|\bm{\mu}-\widehat{\mathbf{X}}_{0,\mathbf{p}_{i}}\|_{2}^{2},$ (11)
@@
With the optimized clean image expectation term $\widetilde{\bm{\mu}}_{t,\mathbf{p}_{i}}$, we further utilize two ways to guide the reverse sampling of the pre-trained SVD: (1) Directly Guided Sampling, which is fast with limited quality; and (2) Posterior Sampling, which is slow but more effective. Specifically, the former directly replaces the $\bm{\mu}_{t,\mathbf{p}}$ in 7 with $\widetilde{\bm{\mu}}_{t,\mathbf{p}_{i}}$ as
@@ 4.2 Adaptive Determination of $\lambda(t,\mathbf{p}_{i})$
Although we have reformulated the NVS-oriented diffusion process in the preceding section, the value of $\lambda(t,\mathbf{p}_{i})$ in Eq. (11) and Eq. (12), which is crucial to the quality of synthesized views, has to be appropriately determined. Here, we theoretically explore the formulation of $\lambda(t,\mathbf{p}_{i})$ via analyzing the upper boundary of the estimation error of $\widetilde{\bm{\mu}}_{t,\mathbf{p}_{i}}$.
@@
$\widehat{\lambda}(t,\mathbf{p}_{i})=\operatorname*{arg\,min}_{\lambda(t,\mathbf{p}_{i})}\frac{1}{1+\lambda(t,\mathbf{p}_{i})}\mathbb{E}_{\mathbf{X}_{t}\sim\mathcal{P}(\mathbf{X}_{t})}(\mathcal{E}_{D})+\frac{\lambda(t,\mathbf{p}_{i})}{1+\lambda(t,\mathbf{p}_{i})}\mathbb{E}_{\mathbf{X}\sim\mathcal{P}(\mathbf{X})}(\mathcal{E}_{P})+v_{1}|\log(\lambda(t,\mathbf{p}_{i}))|,$ (16)
where $v_{1}>0$ and the last regularization term prevents the weights from being overfitting on the empirically estimated errors. In the following, we will provide the explicit formulations of the two error terms $\mathcal{E}_{D}=\|\bm{\mu}_{t,\mathbf{p}_{i}}-\widetilde{\bm{\mu}}_{t,\mathbf{p}_{i}}^{*}\|_{2}$ and $\mathcal{E}_{P}=\|\widehat{\mathbf{X}}_{0,\mathbf{p}_{i}}-\widetilde{\bm{\mu}}_{t,\mathbf{p}_{i}}^{*}\|_{2}$  [^2] There is also inherent discretization error as investigated by (Lu et al., 2022a; b; Zhang & Chen, 2022). However, we could reduce such error terms by enlarging the number of reversing steps., based on theoretical analyses and experimental observations.
Diffusion Estimation Error $\mathcal{E}_{D}$. This error is caused due to the fact that the diffusion model cannot perfectly estimate $\widetilde{\bm{\mu}}_{t,\mathbf{p}_{i}}^{*}$ from the given noised latent $\mathbf{X}_{t}$, Recent works (Zhang & Chen, 2022; Zheng et al., 2024) indicate that the derivative $S_{\theta}(\mathbf{X}_{t},t)$ of SDE-based diffusion models varies intensely when $\sigma(t)$ is huge. Accordingly, large values of $\sigma$ would potentially lead to large errors. Moreover, we experimentally investigated the correlation of $\|\mathcal{E}_{D}\|_{2}$ with $\sigma(t)$. As demonstrated in in Fig. 2 (a). $\|\mathcal{E}_{D}\|_{2}$ gradually decreases with the diffusion reverse process (decreasing of $\sigma(t)$). Thus, we empirically formulate $\|\mathcal{E}_{D}\|_{2}=v_{2}\sigma(t)$, where $0<v_{2}$.
@@
$\widehat{\lambda}(t,\mathbf{p}_{i})=\operatorname*{arg\,min}_{\lambda(t,\mathbf{p}_{i})}\frac{v_{2}\sigma(t)}{1+\lambda(t,\mathbf{p}_{i})}+\frac{\lambda(t,\mathbf{p}_{i})v_{3}\|\Delta\mathbf{p}\|_{2}}{1+\lambda(t,\mathbf{p}_{i})}+v_{1}|\log(\lambda(t,\mathbf{p}_{i}))|,$ (17)
whose closed-form solution is (we have visualized $\widehat{\lambda}$ in Appendix D)
@@
Claim of Novelty. Our method presents a theoretical analysis that links NVS with diffusion processes. Examining the relationships between different views and assessing their potential error distributions enable adaptive modulation of the diffusion score function, effectively minimizing errors from both the warping operator and the diffusion process. This makes our method uniquely suited to addressing NVS challenges. In contrast, traditional guided sampling algorithms typically rely on image degradation models, such as blurring and noise, to guide the diffusion process, setting our approach apart. Moreover, we have theoretically discussed that the proposed method regularizes the diffusion trajectories towards a more accurate direction in Appendix A.1. We have also proved the proposed method (DGS) can safely regularize the samples on the data manifolds in Appendix A.2. We also experimentally compare with inpainting-based methods in Appendix G.